OpenStudy (anonymous):

What is the best way to describe the solution of this system of equation 2x-5y = 8 5y= 2x +8

OpenStudy (anonymous):

2x-5y=8 (then 5y=2x+8) ==> 2x-(2x+8)=8 ==> 2x-2x-8=8 ==>but -8 <>8 this equation is correct.